OpenStudy (kianna0098):

Simplify. (3x^2 + 2x + 7x^3) + (10x^2 + 6x + 9)

OpenStudy (.mehek.):

Add like terms \(3x^2+10x^2=?\\2x+6x\) \(7x^3\) and \(9\) don't have any like terms so leave them as is

OpenStudy (kianna0098):

A: 7x^3+13^2+8+9 B: crossed out C: 13^2+8x+16^3 D: 13x^4+7x^3+8+^2+9

OpenStudy (kianna0098):

@grimesssz

OpenStudy (kianna0098):

what do you think

OpenStudy (.mehek.):

She can't give direct answers. A mod already came and deleted her answer.

OpenStudy (mathstudent55):

\((3x^2 + 2x + 7x^3) + (10x^2 + 6x + 9) =\) Since you are adding polynomials, you don;t need the parentheses, so just drop the parentheses and write: \(=3x^2 + 2x + 7x^3 + 10x^2 + 6x + 9\) Now notice which terms are like terms and move them together. We can do this because of the commutative and associative properties of addition. \(=\color{red}{3x^2} + \color{blue}{2x} + \color{green}{7x^3} + \color{red}{10x^2} + \color{blue}{6x} + 9\) \( =\color{green}{7x^3} + \color{red}{3x^2} + \color{red}{10x^2} +\color{blue}{2x} + \color{blue}{6x} + 9 \) Now combine like terms to find your final answer. Remember that like terms have the same variable part and are shown above in the same color.